OpenStudy (anonymous):

I am a little confused......I am supposed to find GCF of -4, -10 and 20. 2 if the GCF, but is it positive of negative?

OpenStudy (tkhunny):

"Prime" Factorization -4 = -1*2*2 -10 = -1*2*5 20 = 2*2*5